# Unlocking True Elasticity for the Cloud-Native Era with Dandelion

### v1

@@
Abstract
Elasticity is fundamental to cloud computing, as it enables quickly allocating resources to match the demand of each workload as it arrives, rather than pre-provisioning resources to meet performance objectives. However, even serverless platforms — which boot sandboxes in 10s to 100s of milliseconds — are not sufficiently elastic to avoid over-provisioning expensive resources. Today’s FaaS platforms rely on pre-provisioning many idle sandboxes in memory to reduce the occurrence of slow, cold starts. A key obstacle for high elasticity is booting a guest OS and configuring features like networking in sandboxes, which are required to expose an isolated POSIX-like interface to user functions.
Our key insight is that redesigning the interface for applications in the cloud-native era enables co-designing a much more efficient and elastic execution system. Now is a good time to rethink cloud abstractions as developers are building applications to be cloud-native. Cloud-native applications typically consist of user-provided compute logic interacting with cloud services (for storage, AI inference, query processing, etc) exposed over REST APIs. Hence, we propose Dandelion, an elastic cloud platform with a declarative programming model that expresses applications as DAGs of pure compute functions and higher-level communication functions. Dandelion can securely execute untrusted user compute functions in lightweight sandboxes that cold start in hundreds of microseconds, since pure functions do not rely on extra software environments such as a guest OS. Dandelion makes it practical to boot a sandbox on-demand for each request, decreasing performance variability by two to three orders of magnitude compared to Firecracker and reducing committed memory by 96% on average when running the Azure Functions trace.
@@ 1. Introduction
However, even serverless cloud platforms are not elastic enough to avoid pre-provisioning expensive resources. To achieve low latency and high throughput, FaaS platforms keep many idle function sandboxes ready in memory (firecracker, ; serverless-in-the-wild, ). For example, when running the Azure Functions trace (serverless-in-the-wild, ), the Knative (knative, ) serverless platform aggressively autoscales sandboxes such that 97% of requests execute in sandboxes whose memory was provisioned before the invocation arrived (warm starts). Although this improves median latency, it causes the platform to commit 16$\times$ more memory on average (red line in Figure 1) than the memory required for running active requests (blue line). Provisioning so many extra sandboxes in memory is expensive, since DRAM dominates server costs (tmo-dram-cost, ; pond-dram-cost, ).
@@
We show that a key barrier to achieving true elasticity in the cloud is executing each function in a sandbox derived from a container or VM that exposes a POSIX-like interface to the function. A minimal KVM (kvm, ) instance can boot in less than 140 thousand cycles (virtines, ) (i.e. $\sim 52\mu s$ at 2.69GHz). However, booting a hello-world function in a Firecracker MicroVM (which is based on KVM) takes over 10ms even with snapshots. Much of this time is spent on loading the guest OS snapshot and configuring the MicroVM’s network setup (¿8ms on the snapshot demand paging and guest-host connection re-establishment), which are needed to provide a POSIX-like interface to the user function.
Although prioritizing compatibility was essential in the early days of cloud computing to ease migration from on-premise deployments, cloud-native application development is increasingly common today. Developers build applications for cloud environments, leveraging numerous cloud services for storage, analytics, AI inference, and more, which are all exposed over REST APIs (aws-rest-api, ). Cloud-native programming is evolving (serverless-end-linux, ; new-directions-cloud, ; rethinking-serverless, ; compute-centric-networking, ; func_as_func, ; mosaic, ; pulumi, ; yu2023following, ; cloudburst, ; boxer-elasticity, ; serverless-uphill, ; lambdata, ; dataflower, ), with systems like SigmaOS (sigmaos, ) demonstrating that well-tailored interfaces for cloud-native applications can eliminate the need for traditional POSIX-like interfaces. Even today’s FaaS platforms, such as AWS Lambda, are not fully POSIX compliant, as functions cannot accept incoming connections (wawrzoniak2021boxer, ). Hence, now is the right time to rethink cloud abstractions that can enable the underlying infrastructure to become more efficient and truly elastic.
@@
We propose Dandelion, an elastic cloud platform with a declarative cloud-native programming model that replaces the POSIX-based network interface with a higher-level (e.g., HTTP-based) interface for applications to interact with remote services like cloud storage, databases, and AI inference services. Dandelion executes applications expressed as DAGs of pure compute functions and communication functions. Pure compute functions are bodies of (untrusted) user code, which take a declared list of inputs and produce a declared list of outputs. Communication functions are implemented by the Dandelion platform and can be invoked (but not modified) in user application DAGs. Dandelion currently implements one communication function for HTTP communication, which enables REST API interactions, but other protocols can be added. Decomposing a cloud application into pure compute and communication functions enables redesigning the function execution system for high elasticity. First, Dandelion can securely execute pure compute functions in lightweight sandboxes that are fast to cold start, since the sandboxes do not need a guest OS setup for networking, threading, etc. or an isolated virtual network device on the host. Second, Dandelion can dedicate a CPU core for each compute function and run to completion to minimize context switches, since pure functions do not block. [^1] Compute functions do not issue syscalls. Dandelion allocates an isolated memory region for each function before it starts executing, loads inputs in this region (e.g., via a memory-mapped file system), and provides a standard C/C++ library that allows file read/writes without syscalls (see §4.1). Finally, it enables Dandelion to run a cooperative runtime for network I/O per node (rather than per sandbox) and elastically adjust the CPU core allocation based on the number of compute vs. communication functions in the system.
@@ 2.1. FaaS Execution Environment
Traditional VMs can take tens of seconds to boot (boxer-elasticity, ), but MicroVMs on Firecracker (firecracker, ) reduce boot times to hundreds of milliseconds by removing many general-purpose VMM features that FaaS does not need, such as a BIOS and various device emulators. To further reduce boot time to tens of milliseconds (catalyzer, ; reap, ), MicroVMs can be restored from snapshots. Unikernel-based VMs (seuss, ; unikraft, ; hand:unikernels, ; lightvm, ) reduce boot time to the millisecond range by creating a specialized, single-address-space machine image of the user function with only its necessary kernel functionality.
Some cloud or edge platforms run functions as Wasm modules (cloudflare-wasm, ; fastly-wasm, ; faasm, ; sledge, ; spin, ). Wasm is a binary instruction format designed for a stack-based VM to run code at near-native speed with memory safety guarantees provided by a compiler and runtime (wasm, ). Wasm sandboxes boot much faster than MicroVMs, which makes it practical to create sandboxes on-demand per request. However, Wasm code does not necessarily run as fast as native code (lightweight-fault-wasm, ; understanding-wasm-perf, ; sledge, ). Current Wasm compilers may lack optimizations implemented in more mature compilers like Clang, which can result in more instructions, CPU register pressure, and extra branches (not-so-fast-wasm, ), though this is an active area of research and engineering. In our experiments with Wasmtime (wasmtime, ) in §7.3 and §7.6, we found it still performs worse at compute-intensive tasks than native code.
@@ 2.3. Why FaaS is not Elastic Today
Although MicroVM snapshots reduce sandbox boot times down to tens of milliseconds, this latency is still too high to incur for every function invocation. Production traces reveal that many FaaS functions execute for tens of milliseconds or less (huawei-function-trace, ; datadog:state21, ). It is difficult to further reduce this boot latency, as we find that at least 8ms are spent on loading the snapshot (primarily the guest OS state) by demand paging and re-establishing the network connection between the host and the guest process (to provide function inputs and retrieve its outputs). Both are necessary in a POSIX-like sandbox environment.
Wasm sandboxes boot much faster than MicroVMs. However, in practice several commercial Wasm-based serverless platforms still apply defense-in-depth strategies for stronger isolation (though some components of the Wasm toolchain have been formally verified (veriwasm, ; Johnson:WaVe:2023, ; rwasm, )). For example, Cloudflare runs multiple instances of their Wasm runtime as separate processes on each machine to separate low-trusted Wasm workers from more trusted ones. They also use Linux namespaces and seccomp for additional sandboxing at the process level (cloudflare-workers-security, ). This limits elasticity, as scaling now involves starting a Wasm runtime (i.e. loading and initialization overheads) as well as setting up a namespace and seccomp filters.
To give the illusion of microsecond-scale elasticity, cloud providers heavily rely on pre-provisioning idle function sandboxes in memory (serverless-end-linux, ). This has high memory overhead, as shown in Figure 1. Running a guest OS inside each function sandbox also adds to the memory footprint (medes, ; runD, ). Memory usage contributes significantly to the overall cost of operating a FaaS platform, as DRAM dominates server costs (pond-dram-cost, ; tmo-dram-cost, ). This ultimately leads to higher cost for users. Since it is usually prohibitively expensive for providers to keep enough sandboxes hot in memory for all requests (firecracker, ; serverless-in-the-wild, ; faascache, ; knative, ), some requests incur sandbox creation on the critical path, introducing high performance variability for user applications, as shown in Figure 2.
@@ 3. Dandelion’s Approach
To achieve high elasticity while eliminating the need to over-provision idle sandboxes in memory, we propose Dandelion. Our goal is to enable fine-grain (granular-computing, ), elastic computing with secure isolation for a variety of cloud-native applications, such as data transformation pipelines (e.g., image compression, video transcoding (opencv, ; excamera, )), distributed log processing (minio-log-processing, ), CI/CD tasks, elastic query processing (which can include untrusted user-defined function (UDF) execution (containerized-udf, )), and emerging agentic AI workflows (compound-ai-blog, ; agentic-ai-databricks, ).
@@
Our key insight is that redesigning the programming interface for cloud applications in the cloud-native era enables co-designing a more efficient and elastic execution system. We observe that cloud-native applications typically consist of interactions between various cloud services (exposed over REST APIs) and snippets of custom user-provided compute logic. Hence, we propose a programming model in which applications are expressed as direct acyclic graphs (DAGs) of pure compute functions and communication functions. Pure compute functions are snippets of untrusted user code that consume a declared set of inputs and produce a declared set of outputs. Communication functions enable interaction with external services and are implemented by the Dandelion platform. They are invoked from user application DAGs, but for security, their implementation cannot be modified by users. Currently, Dandelion supports an HTTP communication function that enables interaction with remote services like cloud storage buckets, AI inference services, and data warehouses over REST APIs. As we will describe in §4, users provide Dandelion with a DAG of compute and communication functions (which specifies each function’s inputs and outputs) and they also provide the binaries for their custom compute functions. This programming model is inspired by prior dataflow systems (spark, ; beam, ; naiad, ; dryad, ; dryadLINQ, ; malte-dataflow, ). Similar to systems like Spark (spark, ), Dandelion’s programming model makes data dependencies and thus data-parallelism and task-parallelism explicit to the platform. In contrast with traditional dataflow graphs, where edges encode internal communication and external communication is only performed at the dataflow sources and sinks, Dandelion function instantiations can either perform local compute or exchange data with other cloud services.
@@
Decomposing a cloud application into pure compute and communication functions allows Dandelion to optimize the execution system for high elasticity (§5). It enables faster cold starts in Dandelion. Pure compute functions can be securely executed in lightweight sandboxes that provide memory isolation (not necessarily VMs), since they do not rely on extra software environments such as a guest OS. The separation of compute and communication additionally enables Dandelion to use cooperative network I/O and elastically adjust the CPU core allocation based on the number of compute vs. communication functions in the system. The platform can monitor each application’s I/O intensity based on its number of compute vs. communication functions and leverage this information for CPU core allocation, over-committing CPU cores for communication functions and dedicating CPU cores to compute functions.
@@
Non-goals. We design Dandelion for workloads with spiky load patterns (scaling from 0/few to 1/many with long periods of no/low load) or data-dependent parallelism. Many applications can benefit from Dandelion’s ability to quickly scale-out to multiple CPU cores when an application computes on many independent inputs and quickly scale-in when the application needs to process data sequentially. However, Dandelion is not designed for applications that require large, frequent state synchronization, like online transaction processing, online gaming services, AI training, or certain data processing algorithms that rely on fine grained multithreading and shared memory for performance.
@@ 4.1. User Interface
Dandelion provides a composition language to help users express DAGs of compute functions and communication functions in a more developer-friendly syntax. This DSL is inspired by the domain-specific languages that other dataflow systems like Spark (spark, ) and Timely (naiad, ) provide for users to specify DAGs. Our composition language could be replaced with other DSLs. Currently, users need to provide the composition and all compute function binaries directly, but ideally applications could be automatically decomposed into pure compute and communication functions, as we discuss in §4.3.
Communication Functions   Since services generally expose REST APIs (restful_2013, ; rest-api, ), which use HTTP requests to communicate, we have found it sufficient to implement a HTTP communication function. In contrast to compute functions, the communication function contains trusted code that has permissions to invoke system calls. The function performs HTTP protocol-tailored user input sanitization. Support for additional protocols can be added as needed, such as text protocols to communicate with in-memory key-value stores directly over TCP or UDP (memcached-protocol, ).
Compute Functions   Compute functions can use familiar interfaces for memory allocation, local filesystem operations, and basic utilities like math functions, formatting, etc by linking to Dandelion’s custom libc (hlibc) or libc++ (hlibc++). These libraries provide a high-level interface with a userspace in-memory virtual filesystem. The in-memory filesystem represents function input sets and output sets as folders, with items as files within these folders. This enables a compute function to read inputs and write outputs as standard file operations without invoking system calls. Functions requiring system calls (e.g., mmap, mprotect, socket or threading) have stub implementations, returning appropriate error codes. Our current prototype provides SDKs for developing compute functions in C and C++, as well as a Python interpreter (see §4.2).
hlibc and hlibc++ run on top of a lower-level system interface. This interface uses a special data structure that Dandelion sets up before the function starts to point to descriptors for the input sets and items in the function’s memory. Before the function exits, it sets up the structure with the output sets and item descriptors. When a compute function uses the virtual file system to create outputs, hlibc automatically adds all files in folders that are output sets as output items.

### v2

@@
Abstract.
Elasticity is fundamental to cloud computing. An elastic platform can quickly allocate resources to match the demand of each workload as it arrives, rather than pre-provisioning resources to meet performance objectives. However, even serverless platforms — which boot sandboxes in 10s to 100s of milliseconds — are not sufficiently elastic to avoid pre-provisioning expensive resources. Today’s FaaS platforms provision many extra, idle sandboxes in memory to reduce the occurrence of slow, cold starts. Initializing securely isolated sandboxes with a POSIX-like computing environment that today’s cloud users expect is slow as it requires booting a guest OS and configuring networking.
Our key insight is that the rise of cloud-native application development provides an opportunity to rethink the application interface to the cloud and co-design a much more efficient, elastic computing platform under the hood. We propose Dandelion, an elastic cloud platform with a declarative cloud-native programming model that replaces POSIX-based network interfaces with higher-level (e.g., HTTP-based) interfaces for applications to interact with remote services like cloud storage, databases, and AI inference services. Dandelion executes applications expressed as DAGs of pure compute functions and communication functions.
This enables Dandelion to securely execute compute functions in lightweight sandboxes that cold start in 100s of microseconds, since pure functions do not rely on software environments such as a guest OS. Dandelion makes it practical to boot sandboxes on-demand per request, decreasing performance variability by two to three orders of magnitude compared to Firecracker and reducing committed memory by 96% on average when running the Azure Functions trace.
journalyear: 2025copyright: ccconference: ACM SIGOPS 31st Symposium on Operating Systems Principles; October 13–16, 2025; Seoul, Republic of Koreabooktitle: ACM SIGOPS 31st Symposium on Operating Systems Principles (SOSP ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3731569.3764803isbn: 979-8-4007-1870-0/2025/10ccs: Computer systems organization Cloud computing
@@ 1. Introduction
[…]
@@
We show that a key barrier to achieving true elasticity in the cloud is executing each function in a sandbox derived from a container or VM that exposes a POSIX-like interface to the function. A minimal KVM (kvm, ) instance can boot in less than 140 thousand cycles (virtines, ) (i.e. $\sim 52\mu s$ at 2.69GHz). However, booting a hello-world function in a Firecracker MicroVM (which is based on KVM) takes over 10ms even with snapshots. Much of this time is spent on loading the guest OS snapshot and configuring the MicroVM’s network setup (¿8ms is spent on demand paging the snapshot and re-establishing the guest-host connection), which are needed to provide a POSIX-like interface to the user function.
Although prioritizing POSIX compatibility was essential in the early days of cloud computing to ease migration from on-premise deployments, cloud-native application development is increasingly common today. Developers build applications for cloud environments, leveraging numerous cloud services for storage, analytics, AI inference, and more, which are all exposed over REST APIs (aws-rest-api, ). Cloud-native programming is evolving (serverless-end-linux, ; new-directions-cloud, ; rethinking-serverless, ; compute-centric-networking, ; func_as_func, ; mosaic, ; pulumi, ; yu2023following, ; cloudburst, ; boxer-elasticity, ; serverless-uphill, ; lambdata, ; dataflower, ), with systems like SigmaOS (sigmaos, ) demonstrating that well-tailored interfaces for cloud-native applications can eliminate the need for traditional POSIX-like interfaces. Even today’s FaaS platforms, such as AWS Lambda, are not fully POSIX compliant, as functions cannot accept incoming connections (wawrzoniak2021boxer, ). Hence, now is the right time to rethink cloud abstractions that can enable the underlying infrastructure to become more efficient and truly elastic.
@@
We propose Dandelion, an elastic cloud platform with a declarative cloud-native programming model that replaces the POSIX-based network interface with a higher-level (e.g., HTTP-based) interface for applications to interact with remote services like cloud storage, databases, and AI inference services. Dandelion executes applications expressed as direct acyclic graphs (DAGs) of pure compute functions and communication functions. Pure compute functions are bodies of (untrusted) user code, which take a declared list of inputs and produce a declared list of outputs. Communication functions are implemented by the Dandelion platform and can be invoked (but not modified) in user application DAGs. Dandelion currently implements a communication function for HTTP communication, which enables REST API interactions. We plan to add more communication functions to support additional protocols.
Decomposing a cloud application into pure compute and communication functions enables redesigning the function execution system for high elasticity. First, Dandelion can securely execute pure compute functions in lightweight sandboxes that are fast to cold start, since the sandboxes do not need a guest OS setup for networking, threading, etc. or an isolated virtual network device on the host. Second, Dandelion can dedicate a CPU core for each compute function and run it to completion to minimize context switches, since a pure function does not block. [^1] Compute functions do not issue syscalls. Dandelion allocates an isolated memory region for each function before it starts executing, loads inputs in this region (e.g., via a memory-mapped file system), and provides a standard C/C++ library that allows file read/writes without syscalls (see §4.1). Finally, it enables Dandelion to run a cooperative runtime for network I/O per node (rather than per sandbox) and elastically adjust the CPU core allocation based on the number of compute vs. communication functions in the system.
@@ 2.1. FaaS Execution Environment
Traditional VMs can take tens of seconds to boot (boxer-elasticity, ). Firecracker MicroVMs (firecracker, ) reduce boot times to hundreds of milliseconds by removing many general-purpose VMM features that FaaS does not need, such as a BIOS and various device emulators. To further reduce boot time to tens of milliseconds (catalyzer, ; reap, ), MicroVMs can be restored from snapshots. Unikernel-based VMs (seuss, ; unikraft, ; hand:unikernels, ; lightvm, ) reduce boot time to the millisecond range by creating a specialized, single-address-space machine image of the user function with only its necessary kernel functionality.
Some cloud or edge platforms run functions as Wasm modules (cloudflare-wasm, ; fastly-wasm, ; faasm, ; sledge, ; spin, ; hyperlight-wasm, ). Wasm is a binary instruction format designed for a stack-based VM to run code at near-native speed with memory safety guarantees provided by a compiler and runtime (wasm, ). Wasm sandboxes boot much faster than MicroVMs, which makes it practical to create sandboxes on-demand per request. However, Wasm code does not necessarily run as fast as native code (lightweight-fault-wasm, ; understanding-wasm-perf, ; sledge, ). Current Wasm compilers may lack optimizations implemented in more mature compilers like Clang, which can result in more instructions, CPU register pressure, and extra branches (not-so-fast-wasm, ), though this is an active area of research and engineering. Our experiments in §7.3 and §7.6 show that Wasmtime (wasmtime, ) runs slower than native code for compute-intensive tasks.
@@ 2.3. Why FaaS is not Elastic Today
Although MicroVM snapshots reduce sandbox boot times down to tens of milliseconds, this latency is still too high to incur for every function invocation. Production traces reveal that many FaaS functions execute for tens of milliseconds or less (huawei-function-trace, ; datadog:state21, ) and would hence experience high overhead. It is difficult to further reduce boot latency, as we find that at least 8ms are spent on loading a minimal snapshot (primarily the guest OS state) by demand paging and on re-establishing the network connection between the host and the guest process (to provide function inputs and retrieve its outputs). Both are necessary in a POSIX-like sandbox environment.
Using Wasm sandboxes is an option to reduce boot times compared to MicroVMs. However, in practice several commercial Wasm-based serverless platforms still apply defense-in-depth strategies for stronger isolation even though some components of the Wasm toolchain have been formally verified (veriwasm, ; Johnson:WaVe:2023, ; rwasm, ). For example, Cloudflare runs multiple instances of their Wasm runtime as separate processes on each machine to separate low-trusted Wasm workers from more trusted ones. They also use Linux namespaces and seccomp for additional sandboxing at the process level (cloudflare-workers-security, ). This limits elasticity, as scaling now involves starting a Wasm runtime (i.e. loading and initialization overheads) as well as setting up a namespace and seccomp filters.
Since quickly initializing a securely isolated POSIX-like environment for cloud user functions is difficult, today’s platforms pre-provision sandboxes to reduce the probability of incurring guest OS and networking initialization overheads on the critical path of requests. This gives the illusion of microsecond-scale elasticity at the expense of consuming extra memory.
Figure 1 compares the committed memory on a server executing a sample of the Azure Functions trace (serverless-in-the-wild, ) with Knative’s default autoscaling algorithm that provisions extra idle sandboxes (red line) versus the memory usage of active sandboxes over time (blue) line. Running a guest OS inside each function sandbox further adds to the memory footprint (medes, ; runD, ). Memory usage contributes significantly to the overall cost of operating a FaaS platform, as DRAM dominates server costs (pond-dram-cost, ; tmo-dram-cost, ). This ultimately leads to higher cost for users. Since it is usually prohibitively expensive for providers to keep enough sandboxes hot in memory for all requests (firecracker, ; serverless-in-the-wild, ; faascache, ; knative, ), some requests incur sandbox creation on the critical path, introducing high performance variability for user applications, as shown in Figure 2.
@@ 3. Dandelion’s Approach
To achieve high elasticity while eliminating the need to pre-provision idle sandboxes in memory, we propose Dandelion. Our key insight is that redesigning the traditional POSIX-based programming interface for cloud applications in the cloud-native era enables co-designing a more efficient and elastic cloud platform. We observe that cloud-native applications typically consist of interactions between various cloud services (exposed over REST APIs) and snippets of user-defined compute logic. Hence, we propose a programming model in which applications are expressed as DAGs of pure compute functions and communication functions. Pure compute functions are snippets of untrusted user code that consume a declared set of inputs and produce a declared set of outputs. Communication functions enable interaction with external services. They are implemented by the Dandelion platform. Users invoke communication functions from user application DAGs, but for security, the implementation of communication functions cannot be modified by users. Currently, Dandelion supports an HTTP communication function that enables interaction with remote services like cloud storage buckets, AI inference services, and data warehouses over REST APIs. Writing a Dandelion application involves: 1) expressing the DAG of compute and communication functions using a domain-specific language (DSL) and 2) providing the binaries for user-defined compute functions. §4 describes the DSL for expressing application DAGs (which also makes function inputs and outputs explicit), language support for developing user-defined compute functions, and the library of currently supported communication functions. This programming model is inspired by prior dataflow systems (spark, ; beam, ; naiad, ; dryad, ; dryadLINQ, ; malte-dataflow, ). Similar to systems like Spark (spark, ), Dandelion’s programming model makes data dependencies and thus data-parallelism and task-parallelism explicit to the platform. In contrast with traditional dataflow graphs, where edges encode internal communication and external communication is only performed at the dataflow sources and sinks, Dandelion function instantiations can either perform local compute or exchange data with other cloud services. Furthermore, we design Dandelion to execute dataflow DAGs with secure isolation in multi-tenant deployments without relying on execution in traditional VMs, which are slow to boot.
@@
Decomposing a cloud application into pure compute and communication functions allows Dandelion to optimize the execution system for high elasticity while maintaining secure isolation guarantees (§5). It enables faster cold starts: pure compute functions can be securely executed in lightweight sandboxes that provide memory isolation (not necessarily VMs), since they do not rely on extra software environments like a guest OS. The separation of compute and communication also enables Dandelion to use cooperative network I/O and elastically adjust the CPU core allocation based on the number of compute vs. communication functions in the system. The platform can monitor each application’s I/O intensity based on its number of compute vs. communication functions. Dandelion leverages this information for CPU core allocation, over-committing CPU cores for communication functions and dedicating CPU cores to compute functions.
Target applications. We design Dandelion for workloads with spiky load patterns (scaling from 0/few to 1/many with long periods of no/low load) or data-dependent parallelism. Many applications can benefit from Dandelion’s ability to quickly scale-out to multiple CPU cores when an application computes on many independent inputs and quickly scale-in when the application needs to process data sequentially. Our goal is to enable fine-grain (granular-computing, ), elastic computing with secure isolation for a variety of cloud-native applications, such as data transformation pipelines (e.g., image compression, video transcoding (opencv, ; excamera, )), distributed log processing (minio-log-processing, ), CI/CD tasks, elastic query processing (which can include untrusted user-defined function (UDF) execution (containerized-udf, )), and emerging agentic AI workflows (compound-ai-blog, ; agentic-ai-databricks, ).
@@
Non-goals. However, Dandelion is not designed for applications that require large, frequent state synchronization, like online transaction processing, online gaming services, AI training, or data processing algorithms that rely on fine-grain multithreading and shared memory for performance.
@@ 4.1. User Interface
Dandelion provides a composition language to help users express DAGs of compute functions and communication functions in a more developer-friendly syntax. Listing 2 shows an example of the log processing example application expressed as a composition in Dandelion’s DSL. Our DSL is inspired by the domain-specific languages that other dataflow systems like Spark (spark, ) and Timely (naiad, ) provide for users to specify DAGs. Our language could be replaced with other DSLs. Currently, users need to provide the composition and all compute function binaries directly, but ideally applications could be automatically decomposed into pure compute and communication functions, as we discuss in §4.3.
Communication Functions.  Platform-provided communication functions enable user applications to interact with other services. They are designed to support high-level protocols with established user interfaces. A well-defined user interface enables the platform to safely validate requests to protect against malicious requests. It also frees the user from worrying about the lower-level implementation. Since communication functions are trusted, they are allowed to call system calls to establish and manage new connections. While Dandelion could support protocols such as gRPC or SCP, our current implementation focuses on HTTP, since cloud services generally expose REST APIs (restful_2013, ; rest-api, ). Dandelion’s HTTP function first sanitizes inputs and then makes the specified GET/PUT/POST/DELETE requests, handing the response back to the composition for further processing.
Compute Functions.  Compute functions can use familiar interfaces for memory allocation, local filesystem operations, and basic utilities like math functions, formatting, etc by linking to Dandelion’s custom libc (dlibc) or libc++ (dlibc++). These libraries provide a high-level interface with a userspace in-memory virtual filesystem. The in-memory filesystem represents function input sets and output sets as folders, with items as files within these folders. This enables a compute function to read inputs and write outputs as standard file operations without invoking system calls. Functions requiring system calls (e.g., mmap, mprotect, socket or threading) have stub implementations, returning appropriate error codes. Our current prototype provides SDKs for developing compute functions in C and C++, as well as a Python interpreter (see §4.2).
dlibc and dlibc++ run on top of a lower-level system interface. This interface uses a special data structure that Dandelion sets up before the function starts to point to descriptors for the input sets and items in the function’s memory. Before the function exits, it sets up the structure with the output sets and item descriptors. When a compute function uses the virtual file system to create outputs, dlibc automatically adds all files inside folders that represent output sets as output items.